# A Cat Is A Cat (Not A Dog!): Unraveling Information Mix-ups in Text-to-Image Encoders through Causal Analysis and Embedding Optimization

### v4

@@ Abstract
This paper analyzes the impact of causal manner in the text encoder of text-to-image (T2I) diffusion models, which can lead to information bias and loss. Previous works have focused on addressing the issues through the denoising process. However, there is no research discussing how text embedding contributes to T2I models, especially when generating more than one object. In this paper, we share a comprehensive analysis of text embedding: i) how text embedding contributes to the generated images and ii) why information gets lost and biases towards the first-mentioned object. Accordingly, we propose a simple but effective text embedding balance optimization method, which is training-free, with an improvement of 90.05% on information balance in stable diffusion. Furthermore, we propose a new automatic evaluation metric that quantifies information loss more accurately than existing methods, achieving 81% concordance with human assessments. This metric effectively measures the presence and accuracy of objects, addressing the limitations of current distribution scores like CLIP’s text-image similarities. The code is available: https://github.com/basiclab/Unraveling-Information-Mix-ups.
@@ 1 Introduction
Due to the significant information loss issues in text-to-image (T2I) models, such as object mixing and missing, we dissect the generative process of the T2I model to pinpoint the origins of these losses. In the text embedding, the causal manner would make the embedding of the $n^{th}$ token mixed with the token embedding between 00 to $(n-1)^{th}$, which makes the later token embedding similar to the earlier embedding. In the case of a prompt "a cat and a dog", the causal manner would mix the <dog> embedding with the <cat> embedding. This similarity in embeddings results in similar distributions on cross-attention maps, as detailed in Sec.4.4. When denoising directions on these maps align too closely throughout the denoising steps, it can cause a mixed representation if responses are equally low, or one object may overshadow the other if its map elicits a stronger response, leading to object disappearance. These phenomena are visualized in Fig. 1. To address the issue of information bias and loss, we propose the Text Embedding Balance Optimization (TEBOpt) to promote distinctiveness between embeddings of equally important objects for preventing mixing and working alongside existing image latent optimization techniques to address object disappearance. The main contributions of this paper are outlined as follows:
@@
  • We propose the Text Embedding Balance Optimization solution containing one positive and one negative loss to optimize text embedding for tackling information bias with 90.05% improvement in stable diffusion.
@@ 2 Preliminaries
where $\ell_{Q}$, $\ell_{K}$, and $\ell_{V}$ are learned linear projections, and $h_{n}$ represents the head number. $d_{k}$ is the dimension of $\boldsymbol{K}$ and $\mathbf{M}$ is the causal mask. Upon receiving the attention weight $\boldsymbol{Q}\boldsymbol{K}^{\top}\in\mathbb{R}^{{N}\times{N}}$, the mechanism applies the causal mask $\mathbf{M}$. Then, taking $\boldsymbol{V}$ as a weight metric to get the weighted sum of attentions, representing the embedding for each token. Due to the causal manner, the embedding of the $n^{th}$ token ($\varepsilon_{n}$) contains only weighted information for the tokens 00 to $n-1$. This representation causes two issues: First, the earlier token information accumulates the most since every subsequent token has it. Furthermore, the later token’s embedding is not a pure representation of itself, potentially causing identity loss. For example, given a text prompt "a cat and a dog", the earlier token <cat>’s information would be mixed into the embedding of the later token <dog>. It makes the generated image have a higher possibility of generating two cats, as the middle row of Fig. 2 shows.
@@ 3.1.2 De-causal effect for generating images with more than one object
Replacing the token embedding of later mentioned object from the corresponding pure embedding, the hypothesis expects to solve the balance problem and information loss.
@@
Following the setting of the prompt structure (a) in Sec. 3.1, with a text prompt "a/an <obj1> and a/an <obj2>", the token index of <obj1> and <obj2> are taken into the critical token set $O=\{2,5\}$. It first generates the pure embedding of the later mentioned one, "a photo of a <obj2>" ($\varepsilon_{p_{obj2}}$). Then, replace the corresponding embedding dimension of <obj2> in the original embedding ($\varepsilon$) with the pure one. The combined text embedding ($\varepsilon^{\prime}$) is as follows:
@@ 3.2 Method: balancing critical information in the text embedding
While the causal manner results in information bias and information loss, it contributes to generate image content aligned with the general prompt information. To eliminate the accumulated information but retain general information, we design a Text Embedding Balancing Optimization, called TEBOpt (uppermost in Fig. 2), cooperating with image latent optimization to address these issues. The TEBOpt tackles information bias and object mixture issues while the latent optimization tackles object mixture and missing issues. Since object missing is caused by insufficient response value and region activated in the cross-attention map corresponding to the missing object, text embedding cannot specify the activated position; therefore, our method cooperates with existing latent optimization methods to tackle the issue.
@@ 4.3 Qualitative results
Fig. 4 demonstrated the visual comparison between all methods. The same color of the bounding box and the underline for the prompt indicate the same object. The bounding boxes are predicted based on our proposed evaluation method discussed in Sec. 4.2. As an object is wrapped by two bounding boxes with different colors, e.g., A&E [1] in the first row or Structure Diffusion [3] in the second row, it has a high potential to contain mixed objects. In the first row in Fig. 4, our TEBOpt ($\mathcal{L}_{TEB}$) helps Stable Diffusion generate the specified dog and help A&E to make the horse dissimilar from the dog. In the second row, our TEBOpt ($\mathcal{L}_{TEB}$) helps generate a cat less similar to a lion, which is initially equipped with a mane as a visual signal of a lion. Regarding Fig. 5, we demonstrate the generated image with and without our proposed TEBOpt ($\mathcal{L}_{TEB}$). Using our TEBOpt, T2I models solved both the object mixture and missing issues, especially when trying to solve the problem of object mixtures. It is worth noting that objects mixture or missing are also affected by the denoising process in T2I models. Our text optimization mainly contributes to balancing the information in the text embedding and further reducing mixed and missing issue.
@@ 4.4 Quantitative results
Prompt default +$\mathcal{L}_{TEB}$
Text sim 0.454 0.286
   max 0.669 0.537
   min 0.291 0.137
Map dist 2.321 2.637
   max 5.722 6.340
   min 0.287 0.553
Table 5: Analysis of optimized token embedding similarity and cross-attention map distance between two objects within one prompt.
Table 4.4 evaluates the similarity of token embeddings and the distance of cross-attention maps between two objects within one prompt. Token embedding similarity (Text sim) is calculated by cosine similarity while the cross-attention map distance (Map dist) is calculated by the symmetric Kullback-Leibler divergence between two normalized cross-attention maps $M_{i}$ and $M_{j}$: $\frac{1}{2}(D_{KL}(M_{i}||M_{j})+D_{KL}(M_{j}||M_{i}))$, where $D_{KL}(M_{i}||M_{j})=\sum_{pixels}M_{i}\log(M_{i}/M_{j})$. With our $\mathcal{L}_{TEB}$, the token embedding similarity between two objects reduces 36.98% and the cross-attention map distance increases 13.62%.
@@
We calculated the cosine similarity between various text embeddings and displayed the results in Fig. 6 (a). The data indicates that objects with similar colors or sizes, e.g., penguin-panda or turtle-frog, tend to exhibit higher similarity in their text embeddings, which can be attributed to the training mechanism of CLIP. Additionally, we computed the distance between the cross-attention maps, using the same function in Table 4.4, generated by the two objects’ text embeddings with the same initial latent in the early denoising steps. As shown in Fig. 6 (b), there is a positive correlation between the similarity of text embeddings and the distance of the cross-attention maps triggered by the two objects. Specifically, objects with similar text embeddings are more likely to activate overlapping areas during the denoising process. This confirms that similar text embeddings contribute to object mixture, while the short distance of cross-attention maps leading to object missing has been proven by SynGen [17].
@@ 6 Conclusion
In this study, we conducted a detailed analysis of text embedding’s impact on text-to-image diffusion models, a topic rarely explored. Our findings indicate that the causal processing of text embedding leads to information accumulation, causing biases and loss. Directly replacing accumulated embeddings with purified embeddings, though resulting in decreased coexistence of two objects, enhances the balance between generating either object. We introduce a training-free Text Embedding Balance Optimization (TEBOpt) method that effectively eliminates problematic information in key token embeddings, improving information balance handling in stable diffusion by 90.05% while preserving dual-object performance. Additionally, due to the unreliability of existing metrics for assessing inaccuracies in generated images, we propose a new automatic evaluation metric to more effectively measure information loss.
@@ Appendix B Implementation details
All the analyses are conducted on Stable Diffusion 1.4, where the text encoder is the pre-trained CLIP ViT-L/14 [16] and the scheduler is the PNDMSchedular. In the text encoder CLIP, the maximum token length (N) is 77, the embedding dimension (D) is 768, and the attention head number ($h_{n}$) is 12. We use a fixed guidance scale of 7.5 and set the denoising step to 50. For performing the TEB loss in Equation 6, we set a maximum optimization time as 20 and a threshold for $\mathcal{L}_{TEB}$ as -0.7, which comes from $\mathcal{L}^{pos}_{TEB}=0.95$ and $\mathcal{L}^{neg}_{TEB}=0.25$. The optimization target for $\mathcal{L}_{TEB}$ is -0.7 but if the optimization time exceeds 20, it will stop optimizing the text embedding. Moreover, all baselines are conducted with their official source codes.
@@ Appendix C Existing evaluation methods discussion
C.1 Demonstration of bounding boxes interaction corresponding to the evaluation status.
Considering the working mechanism of SOTA text-to-image models, when the cross-attention maps of 2 objects respond closely during denoising, there is a high probability of generating a mixture object, as Fig. 10 shows. Consequently, within the nature of the Owl-ViT detector, these mixtures can be identified by their high overlap.
@@ Appendix D Analysis of information bias with more than 2 objects
Table 6: Analysis of information bias in multiple objects. When there are more objects in the prompt, the bias might gradually enlarge as the order of the mentioned objects gets farther from the first mentioned object. In this setting, we utilize the same evaluation metric to define the information bias among three objects, which is divided into pairs of (obj1, obj2), (obj1, obj3), and (obj2, obj3). As the information bias approaches 1, it yields more balanced results regarding the existence of the object.
@@ Appendix E More qualitative and quantitative comparisons in T2I-CompBench [6]
More qualitative results are in the Fig. 11. The public benchmark, e.g., T2I-CompBench, contains more diverse prompts, e.g., adjective binding with noun, that would make the experiment distract. Thus, in the main paper, we follow related works, e.g., A&E (SIGGRAPH’23) [1] and Predicated diffusion (CVPR’24) [21], to create the most suitable benchmark for our task. While we respectfully disagree with the directness to conduct public T2I-CompBench to demonstrate our performance, we still conduct quantitative and qualitative comparisons in T2I-CompBench and prove our effectiveness.
@@ E.1 Color set
We include a quantitative comparison in Table 7 to compare SD 1.4, SD 1.5, ELLA [5] (on the only released version, SD 1.5), SDXL-Turbo [20], and SD3 [2] on the color set within the T2I-CompBench [6] with 1,000 cases. Within all the baseline methods, our TEBOpt improves the 2 object co-existence rate with 7.9% on SD3 by addressing object mixture and object missing. Also, TEBOpt makes the generated information more balanced, where the info bias is from 2.07 to 1.30.
@@
Table 7: Quantitative comparison with SOTA methods on the color set in the T2I-CompBench.
@@ E.2 Spatial set
We conducted our TEBOpt experiments using Stable Diffusion (SD) 1.4 on the set of spatial relationships (e.g., "next to," "on the side of," etc.) within the T2I-CompBench [6]. This dataset includes nouns representing 5 types of people (man, girl, etc.), 16 types of animals (giraffe, turtle, etc.), and 30 types of objects (table, car, etc.), encompassing a total of 1,000 cases. Table 8 shows that our method demonstrates improvement in increasing the 2 object co-existence with 6.8% and reducing the information bias from 1.43 to 1.21. In this experiment, we further prove that when two nouns in the given prompt are from different categories, such as "a woman and a chair," resulting in a larger text embedding distance, it causes the mixture issue in text-to-image models to be concealed beneath the surface. Thus, while our paper focuses on the task of handling only animals, we reveal and address both the mixture and missing issue. Furthermore, we demonstrate that our method is effective across a more diverse set of categories.
@@
Figure 11: More qualitative results on SD 1.4, ELLA [5] on SD 1.5, SDXL-Turbo [20], and SD3 [2] in complex prompts from color and spatial sets within T2I-CompBench [6].

### v5

@@ Abstract
This paper analyzes the impact of causal manner in the text encoder of text-to-image (T2I) diffusion models, which can lead to information bias and loss. Previous works have focused on addressing the issues through the denoising process. However, there is no research discussing how text embedding contributes to T2I models, especially when generating more than one object. In this paper, we share a comprehensive analysis of text embedding: i) how text embedding contributes to the generated images and ii) why information gets lost and biases towards the first-mentioned object. Accordingly, we propose a simple but effective text embedding balance optimization method, which is training-free, with an improvement of 125.42% on information balance in stable diffusion. Furthermore, we propose a new automatic evaluation metric that quantifies information loss more accurately than existing methods, achieving 81% concordance with human assessments. This metric effectively measures the presence and accuracy of objects, addressing the limitations of current distribution scores like CLIP’s text-image similarities. The code is available: https://github.com/basiclab/Unraveling-Information-Mix-ups.
@@ 1 Introduction
Due to the significant information loss issues in T2I models, such as object mixing and missing, we dissect the generative process of the T2I model to pinpoint the origins of these losses. In the text embedding, the causal manner would make the embedding of the $n^{th}$ token mixed with the token embedding between 00 to $(n-1)^{th}$, which makes the later token embedding similar to the earlier embedding. In the case of a prompt "a cat and a dog", the causal manner would mix the <dog> embedding with the <cat> embedding. This similarity in embeddings results in similar distributions on cross-attention maps, as detailed in Sec.5. When denoising directions on these maps align too closely throughout the denoising steps, it can cause a mixed representation if responses are equally low, or one object may overshadow the other if its map elicits a stronger response, leading to object disappearance. These phenomena are visualized in Fig. 1. To address the issue of information bias and loss, we propose the Text Embedding Balance Optimization (TEBOpt) to promote distinctiveness between embeddings of equally important objects for preventing mixing and working alongside existing image latent optimization techniques to address object disappearance. The main contributions of this paper are outlined as follows:
@@
  • We propose the Text Embedding Balance Optimization solution containing one positive and one negative loss to optimize text embedding for tackling information bias with 125.42% improvement in stable diffusion.
@@ 2 Preliminaries
where $\ell_{Q}$, $\ell_{K}$, and $\ell_{V}$ are learned linear projections, and $h_{n}$ represents the head number. $d_{k}$ is the dimension of $\boldsymbol{K}$ and $\mathbf{M}$ is the causal mask. Upon receiving the attention weight $\boldsymbol{Q}\boldsymbol{K}^{\top}\in\mathbb{R}^{{N}\times{N}}$, the mechanism applies the causal mask $\mathbf{M}$. Then, taking $\boldsymbol{V}$ as a weight metric to get the weighted sum of attentions, representing the embedding for each token. Due to the causal manner, the embedding of the $n^{th}$ token ($\varepsilon_{n}$) contains weighted information for the tokens 00 to $n-1$. This representation causes two issues: First, the earlier token information accumulates the most since every subsequent token has it. Furthermore, the later token’s embedding is not a pure representation of itself, potentially causing identity loss. For example, given a text prompt "a cat and a dog", the earlier token <cat>’s information would be mixed into the embedding of the later token <dog>. It makes the generated image have a higher possibility of generating two cats, as the middle row of Fig. 2 shows.
@@ 3.1.2 De-causal effect for generating images with more than one object
Replacing the token embedding of later mentioned object from the corresponding pure embedding, the hypothesis expects to solve the bias problem and information loss.
@@
Following the setting of the prompt structure (a) in Sec. 3.1, with a text prompt "a/an <obj1> and a/an <obj2>", the token index of <obj1> and <obj2> are taken into the critical token set $O=\{2,5\}$. It first generates the pure embedding of the later mentioned one, "a photo of a <obj2>" ($\varepsilon_{p_{obj2}}$). Then, replace the corresponding token embedding of <obj2> in the original embedding ($\varepsilon$) with the pure one. The combined text embedding ($\varepsilon^{\prime}$) is as follows:
@@ 3.2 Method: balancing critical information in the text embedding
While the causal manner results in information bias and information loss, it contributes to generate image content aligned with the general prompt information. To eliminate the accumulated information but retain general information, we design a Text Embedding Balancing Optimization, called TEBOpt (uppermost in Fig. 2), cooperating with image latent optimization to address these issues. The TEBOpt tackles information bias and object mixture issues while the latent optimization tackles object mixture and missing issues. Since the missing object is caused by an insufficient response value and an inadequately activated region in the cross-attention map corresponding to that object, and the text embedding is unable to precisely determine the activated position, our method cooperates with existing latent optimization methods to address this issue.
@@ 4.3 Qualitative results
Fig. 4 demonstrated the visual comparison between all methods. The same color of the bounding box and the underline for the prompt indicate the same object. The bounding boxes are predicted based on our proposed evaluation method discussed in Sec. 4.2. As an object is wrapped by two bounding boxes with different colors, e.g., A&E [1] in the first row or Structure Diffusion [3] in the second row, it has a high potential to contain mixed objects. In the first row in Fig. 4, our TEBOpt ($\mathcal{L}_{TEB}$) helps Stable Diffusion generate the specified dog and help A&E to make the horse dissimilar from the dog. In the second row, our TEBOpt ($\mathcal{L}_{TEB}$) helps Stable Diffusion generate a cat less similar to a lion, which is initially equipped with a mane as a visual signal of a lion. Regarding Fig. 5, we demonstrate the generated image with and without our proposed TEBOpt ($\mathcal{L}_{TEB}$). Using our TEBOpt, T2I models solved both the object mixture and missing issues, especially when trying to solve the problem of object mixtures. It is worth noting that objects mixture or missing are also affected by the denoising process in T2I models. Our text optimization mainly contributes to balancing the information in the text embedding and further reducing mixed and missing issue.
@@ 4.4 Quantitative results
In Table 5, we evaluate the similarity of token embeddings and the distance of cross-attention maps between two objects within one prompt. Token embedding similarity (Token sim) is calculated by cosine similarity while the cross-attention map distance (Map dist) is calculated by the symmetric Kullback-Leibler divergence between two normalized cross-attention maps $M_{i}$ and $M_{j}$: $\frac{1}{2}(D_{KL}(M_{i}||M_{j})+D_{KL}(M_{j}||M_{i}))$, where $D_{KL}(M_{i}||M_{j})=\sum_{pixels}M_{i}\log(M_{i}/M_{j})$. With our $\mathcal{L}_{TEB}$, the token embedding similarity between two objects reduces 36.98% and the cross-attention map distance increases 13.62%.
Token sim $\downarrow$ SD 1.4 +$\mathcal{L}_{TEB}$ Map dist $\uparrow$ SD 1.4 +$\mathcal{L}_{TEB}$
Average 0.454 0.286 (-36.98%) Average 2.321 2.637 (+13.62%)
$[\min,\max]$ [0.291, 0.669] [0.137, 0.537] $[\min,\max]$ [0.287, 5.722] [0.553, 6.340]
Table 5: Analysis of optimized token embedding similarity (Token sim) and cross-attention map distance (Map dist) between two objects within one prompt.
@@
We calculated the cosine similarity between various text embeddings and displayed the results in Fig. 6 (a). The data indicates that objects with similar colors or sizes, e.g., penguin-panda or turtle-frog, tend to exhibit higher similarity in their text embeddings, which can be attributed to the training mechanism of CLIP. Additionally, we computed the distance between the cross-attention maps, using the same function in Table 5, generated by the two objects’ text embeddings with the same initial latent in the early denoising steps. As shown in Fig. 6 (b), there is a positive correlation between the similarity of text embeddings and the distance of the cross-attention maps triggered by the two objects. Specifically, objects with similar text embeddings are more likely to activate overlapping areas during the denoising process. This confirms that similar text embeddings contribute to object mixture, while the short distance of cross-attention maps leading to object missing has been proven by SynGen [16].
@@ 6 Conclusion
In this study, we conducted a detailed analysis of text embedding’s impact on text-to-image diffusion models, a topic rarely explored. Our findings indicate that the causal processing of text embedding leads to information accumulation, causing biases and loss. Directly replacing accumulated embeddings with purified embeddings, though resulting in decreased coexistence of two objects, enhances the balance between generating either object. We introduce a training-free Text Embedding Balance Optimization (TEBOpt) method that effectively eliminates problematic information in critical token embeddings, improving information balance handling in stable diffusion by 125.42% while preserving object coexistence performance. Additionally, due to the unreliability of existing metrics for assessing inaccuracies in generated images, we propose a new automatic evaluation metric to more effectively measure information loss.
@@ Appendix B Implementation details
All the analyses are conducted on Stable Diffusion 1.4, where the text encoder is the pre-trained CLIP ViT-L/14 [15] and the scheduler is the PNDMSchedular. In the text encoder, CLIP, the maximum token length (N) is 77, the embedding dimension (D) is 768, and the attention head number ($h_{n}$) is 12. We use a fixed guidance scale of 7.5 and set the denoising step to 50. For performing the TEB loss in Equation 6, we set a maximum optimization time as 20 and a threshold for $\mathcal{L}_{TEB}$ as -0.7, which comes from $\mathcal{L}^{pos}_{TEB}=0.95$ and $\mathcal{L}^{neg}_{TEB}=0.25$. The optimization target for $\mathcal{L}_{TEB}$ is -0.7 but if the optimization time exceeds 20, it will stop optimizing the text embedding. Moreover, all baselines are conducted with their official source codes.
@@ Appendix C Existing evaluation methods discussion
C.1 Demonstration of bounding boxes interaction corresponding to the evaluation status
Considering the working mechanism of SOTA text-to-image models, when the cross-attention maps of 2 objects respond closely during denoising, there is a high probability of generating a mixture object, as Fig. 10 shows. Consequently, within the nature of the Owl-ViT detector, these mixtures can be identified by their high overlap with high confidence.
@@ Appendix D Analysis of information bias with more than 2 objects
Table 6: Analysis of information bias in multiple objects. When there are more objects in the prompt, the bias might gradually enlarge as the order of the mentioned objects gets farther from the first mentioned object. In this setting, we utilize the same evaluation metric to define the information bias among three objects, which is divided into pairs of (obj1, obj2), (obj1, obj3), and (obj2, obj3). As the info bias approaches 1, it yields more balanced results regarding the existence of the object.
@@ Appendix E More qualitative and quantitative comparisons in T2I-CompBench [5]
More qualitative results are in Fig. 11, Fig. 12, Fig. 13, and Fig. 14. The public benchmark, e.g., T2I-CompBench [5], contains more diverse prompts, e.g., adjective binding with noun, that would make the experiment distract. Thus, in the main paper, we follow related works, e.g., A&E (SIGGRAPH’23) [1] and Predicated diffusion (CVPR’24) [20], to create the most suitable benchmark for our task. While we respectfully disagree with the directness to conduct public T2I-CompBench to demonstrate our performance, we still conduct quantitative and qualitative comparisons in T2I-CompBench and prove our effectiveness.
@@ E.1 Color set
We include a quantitative comparison in Table 7 to compare SD 1.4, SD 1.5, ELLA (ArXiv’24; on the only released version, SD 1.5), SDXL-Turbo [19], and SD3 [2] on the color set within the T2I-CompBench [5] with 1,000 cases. Within all the baseline methods, our TEBOpt improves the 2 object co-existence rate with 7.9% on SD3 by addressing object mixture and object missing. Also, TEBOpt makes the generated information more balanced, where the info bias is from 2.07 to 1.30.
@@
Table 7: Quantitative comparison with SOTA methods on the color set in the T2I-CompBench. Reference: ELLA: Equip Diffusion Models with LLM for Enhanced Semantic Alignment (ArXiv’24)
@@ E.2 Spatial set
We conducted our TEBOpt experiments using Stable Diffusion (SD) 1.4 on the set of spatial relationships (e.g., "next to," "on the side of," etc.) within the T2I-CompBench [5]. This dataset includes nouns representing 5 types of people (man, girl, etc.), 16 types of animals (giraffe, turtle, etc.), and 30 types of objects (table, car, etc.), encompassing a total of 1,000 cases. Table 8 shows that our method demonstrates improvement in increasing the 2 object co-existence with 6.8% and reducing the information bias from 1.43 to 1.21. In this experiment, we further prove that when two nouns in the given prompt are from different categories, such as "a woman and a chair," resulting in a larger text embedding distance, it causes the mixture issue in text-to-image models to be concealed beneath the surface. Thus, while our main paper focuses on the task of handling only animals, it contains different challenges compared to handling more diverse objects in different categories. In our focus, we reveal and address both the mixture and missing issue. Furthermore, in the following experiment, we demonstrate that our method is effective across a more diverse set of categories.
@@
Figure 11: More qualitative results on SD 1.4 in complex prompts from color and spatial sets within T2I-CompBench [5].
Figure 12: More qualitative results on ELLA on SD 1.5 in complex prompts from color set within T2I-CompBench [5]. Reference: ELLA: Equip Diffusion Models with LLM for Enhanced Semantic Alignment (ArXiv’24)
Figure 13: More qualitative results on SDXL-Turbo [19] in complex prompts from color set within T2I-CompBench [5].
Figure 14: More qualitative results on SD3 [2] in complex prompts from color set within T2I-CompBench [5].